# VRWKV-Editor: Reducing Quadratic Complexity in Transformer-Based Video Editing

### v2

@@ 2 Related works
This section discusses four research areas related to our approach; video editing methods using diffusion models, which highlight the computational challenges inherent in current approaches; linear attention mechanisms, which offer efficiency gains over quadratic attention; and computational efficiency approaches in video processing, which directly motivate our integration of linear attention for scalable video editing.
@@ 2.1 Video Editing with Diffusion Models
Complementary to these editing approaches, advances in video generation architectures have explored pixel-based, latent-based, and hybrid designs [14, 21, 22]. Research has shown that pixel-based diffusion models excel in generating low-resolution videos with superior text-video alignment and motion fidelity, while latent-based models offer computational efficiency but may compromise alignment quality [22]. Hybrid approaches that combine pixel-based keyframe generation with latent-based super-resolution have demonstrated the potential to achieve both high quality and computational efficiency [23]. Effective video models incorporate temporal layers within U-Net architectures, including temporal convolution layers and temporal attention mechanisms to facilitate dynamic temporal data assimilation [24]. The integration of spatial and temporal processing remains crucial for maintaining coherent motion patterns across video sequences [25].
@@ 2.2 Linear Attention Mechanisms in Vision
The computational limitations of traditional attention mechanisms have sparked interest in linear attention alternatives that maintain performance while reducing complexity [26, 27]. Recent advances in natural language processing have introduced models with linear feature aggregation mechanisms, such as RWKV and Mamba [4, 5], which emerge as efficient solutions for processing lengthy sequences while maintaining transformer-like capabilities. These architectures demonstrate that linear attention can effectively capture long-range dependencies without the quadratic computational overhead of traditional self-attention. The adaptation of RWKV architectures for diffusion models has shown promising results in addressing computational complexity limitations. Diffusion-RWKV [28] demonstrates that RWKV-based architectures can efficiently handle patchified inputs in diffusion models while significantly reducing spatial aggregation complexity, making them exceptionally adept at processing high-resolution images without requiring windowing or group cached operations. This work establishes the feasibility of linear attention mechanisms in generative modeling tasks, achieving performance comparable to transformer-based approaches while reducing computational overhead. VRWKV [8] adapts linear attention principles to vision tasks, demonstrating robust global processing capabilities while efficiently handling sparse inputs and maintaining linear computational complexity. The bidirectional processing capabilities of VRWKV’s Bi-WKV mechanism offer particular promise for tasks requiring comprehensive spatial-temporal understanding (e.g. video editing). Building upon the success of Diffusion-RWKV in image generation, the extension to video editing tasks represents a natural progression toward more efficient video processing architectures.
@@ 3 Background
This section provides the necessary foundations that form the backbone of our work. We first discuss latent variable refinement techniques that improve efficiency while preserving quality. We then review diffusion models in the context of video generation and highlight their extension from image-based frameworks to spatio-temporal domains, followed by a detailed overview of the VRWKV architecture, which offers a linear-complexity alternative to traditional attention mechanisms. These components form the conceptual backbone for our proposed method.
@@ 3.1 Diffusion Models and Latent Variable Refinement
Diffusion models represent [36] a powerful class of generative models that operate through a progressive noise addition and removal process. Starting with an original image $x_{0}$, the forward diffusion process gradually corrupts the data by adding Gaussian noise over $T$ timesteps. At each step $t$, the noisy image $x_{t}$ is obtained through:
@@
The reverse process learns to denoise these corrupted images, effectively modeling the data distribution. For conditional generation tasks, we can extend this framework to model the conditional distribution $p_{\theta}(x_{t}|c,x_{0})$, where $c$ represents conditioning information such as text prompts or class labels. The conditional model learns:
@@
where $x_{T}\sim N(0,I)$ represents pure noise.
The model parameters $\theta$ are optimized by maximizing the variational lower bound of the log-likelihood:
$\mathcal{L}=\mathbb{E}_{x_{0},c,t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|^{2}\right],$ (4)
where the network $\epsilon_{\theta}$ learns to predict the noise added at each timestep given the noisy image $x_{t}$, timestep $t$, and conditioning information $c$.
To improve computational efficiency, latent diffusion models perform the diffusion process in a compressed latent space rather than directly in pixel space. This approach utilizes a pre-trained autoencoder consisting of an encoder $E$ that maps images to latent representations $z=E(x)$ and a decoder $D$ that reconstructs images from latents.
The diffusion process then operates in this latent space, learning the distribution of encoded representations. This approach significantly reduces memory and computational costs while maintaining high-quality generation capabilities [37]. During inference, the model generates latent representations through the reverse diffusion process, which are then decoded back to pixel space using the pre-trained decoder.
@@ 3.2 Diffusion Models for Video Processing
Beyond their success in image generation, diffusion models have recently been extended to video synthesis, where they enable high-quality spatiotemporal generation and editing [38, 39, 40]. These models operate by learning to reverse a gradual noise addition process, where data is progressively corrupted through a forward Markov chain and then restored through a learned reverse process.
The transition from image to video generation involves inflating 2D models to handle spatio-temporal dimensions, typically through pseudo-3D convolutions where 3×3 kernels are replaced by 1×3×3 kernels, and temporal attention layers are added to transformer blocks [22, 41]. However, this extension introduces significant computational challenges due to the quadratic complexity of attention mechanisms when processing multiple frames simultaneously.
@@ 3.3 VRWKV Architecture
Figure 2: VRWKV [8] module employed on VRWKV-Editor. The VRWKV architecture includes identical VRWKV encoder layer, an average pooling layer, and a linear prediction head. Q-Shift denotes the quad-directional shift method tailed for vision tasks.
VRWKV consists of three main components: Receptance $(R)$, Key $(K)$, and Value $(V)$ transformations, combined with a time-mixing mechanism specifically adapted for vision tasks. For a given input sequence of video frames $X_{t}\in\mathbb{R}^{f\times c\times h\times w}$, where $f$ is the number of frames, $c$ is the number of channels, and $(h,w)$ are the spatial dimensions, the VRWKV operations are defined as:
@@
$W_{R}$, $W_{K}$, $W_{V}$ are learnable projection matrices, $\mu_{R}$, $\mu_{K}$, $\mu_{V}$ are learnable interpolation factors, and $\otimes$ denotes a spatiotemporal convolution, defined as a standard $2D$ spatial convolution applied on a temporally interpolated frame representation $\mu\odot X_{t}+(1-\mu)\odot X_{t-1}$. This differs from standard 3D convolution since it only considers the current and previous frame, rather than a full temporal kernel.
@@
$wkv_{t}=\operatorname{Bi-WKV}(K,V)_{t}=\frac{\sum_{i=0,i\neq t}^{T-1}e^{-(|t-i|-1)/T\cdot w+k_{i}}v_{i}+e^{u+k_{t}}v_{t}}{\sum_{i=0,i\neq t}^{T-1}e^{-(|t-i|-1)/T\cdot w+k_{i}}+e^{u+k_{t}}}.$ (8)
where $T$ represents the total number of tokens, $w$ and $u$ are two $d$-dimensional learnable vectors that represent channel-wise spatial decay and the bonus indicating the current token, respectively. $k_{t}$ and $v_{t}$ denotes $t$-th feature of $K$ and $V$.
@@
$Y_{t}=X_{t}+O_{t}+CM_{t},$ (13)
Figure 3: Pipeline of VRWKV-Editor: Given a text–video pair (e.g., “A man with a backpack hikes on a rocky terrain”) as input, our method leverages pretrained text-to-image diffusion models for text-to-video generation. The input video is first encoded into a discrete latent space, after which our U-Net architecture predicts the injected noise. During inference, a novel video is synthesized by inverting the discrete noise from the input video, guided by an edited prompt (e.g., “An astronaut with a jetpack floats above a Martian landscape, with red rocky terrains and tall”).
@@ 4 Method
This section outlines our method, we begin by formulating the problem and highlighting the computational efficiency challenges encountered in existing architectures. This serves as the primary motivation for our approach, we then introduce the proposed architecture.
@@ 4.1 Problem Formulation
Given a video sequence $V=\{v_{i}\}_{1}^{n}$ with $n$ frames, and a text prompt $P$ describing $V$, our goal is to generate an edited video $V^{{}^{\prime}}=\{v_{i}^{{}^{\prime}}\}_{1}^{n}$ that aligns with new prompts $P^{{}^{\prime}}$ provided by the user, where $v_{i}^{{}^{\prime}}\in\mathbb{R}^{C\times H\times W}$ is a frame with C channels, height H and width W.
@@ 4.2 Computational Efficiency Challenges
where Q, K and V represents the Querie, Key and Value respectively. Although effective in capturing long-range dependencies, this operation scales poorly with sequence length.
@@
4.3 Overview of VRWKV-Editor
@@
Tune-A-Video [11] reduced the high computational complexity of spatio-temporal attention (ST-Attn) in generating videos with increasing frames by introducing a novel sparse version of causal attention mechanism, where the attention matrix is computed between frame $z_{v_{i}}$ and two previous frames $z_{v_{1}}$ and $z_{v_{i-1}}$, reducing the complexity from $O((dT)^{2})$ to $O(2dT^{2})$. However, the quadratic complexity still remains a bottleneck for long sequences of video frames, limiting their ability to efficiently process high-resolution images and lengthy sequences, posing a significant barrier to their broader application. To overcome this issue, we propose a linear attention mechanism based on the VRWKV (Vision Receptance Weighted Key Value) framework, where we replace the traditional quadratic attention computation with a receptance-weighted key-value operation that scales linearly with sequence length, reducing the complexity from $O(2dT^{2})$ to $O(2dT)$. Our approach leverages the VRWKV’s inherent ability to capture long-range dependencies through its recurrent formulation while maintaining spatial awareness through learnable shift operations and channel mixing.
@@
An overview of our approach is depicted in Figure 3. Given a text-video pair as input, our model leverages the pretrained T2I diffusion models for T2V generation by encoding the input video into a discrete space, then we predict the added noise using our U-Net architecture enhanced with VRWKV modules. During inference, we sample a novel video from the discrete noise inverted from the input video, guided by an edited prompt.
4.4 Vision-Recurrent Weighted Key-Value (VRWKV) Module
@@
During our experiments, we observed that traditional CNN-based spatiotemporal modules failed to capture complex spatiotemporal dependencies, while adding more cross-attentions or Temporal Attentions only increased the complexity of the model without improving the quality of edited videos or maintaining their consistency. Inspired by the RWKV architecture’s ability to combine the strengths of RNNs and Transformers, we introduce a variant of Vision-Recurrent Weighted Key-Value (VRWKV) module specifically designed for video spatiotemporal modeling. The VRWKV module addresses key limitations of traditional attention mechanisms in video editing. Unlike quadratic attention mechanisms, VRWKV operates with linear complexity $O(d)$, making it more efficient for processing long video sequences (Table 2). The module maintains a recurrent state that captures temporal dependencies across frames while preserving spatial relationships. Moreover, the weighted key-value mechanism allows for dynamic focus on relevant spatiotemporal features.
@@
Figure 4: Qualitative comparison between VRWKV-Editor [44] and CAMEL [45], CCEdit [46], ControlVideo [20], Swin-Editor [44] and Vid2Vid-zero [10]. The first row presents the input video. The figure highlights two different scenarios: Object editing and object-background editing . The results showcase the superiority of our framework in editing videos across various scenarios, producing consistent and high quality outputs.

### v3

@@ 2 Related works
This section discusses three research areas related to our approach; video editing methods using diffusion models, which highlight the computational challenges inherent in current approaches; linear attention mechanisms, which offer efficiency gains over quadratic attention; and computational efficiency approaches in video processing, which directly motivate our integration of linear attention for scalable video editing.
@@ 2.1 Video Editing with Diffusion Models
Complementary to these editing approaches, advances in video generation architectures have explored pixel-based, latent-based, and hybrid designs [14, 21, 22, 23]. Research has shown that pixel-based diffusion models excel in generating low-resolution videos with superior text-video alignment and motion fidelity, while latent-based models offer computational efficiency but may compromise alignment quality [23]. Hybrid approaches that combine pixel-based keyframe generation with latent-based super-resolution have demonstrated the potential to achieve both high quality and computational efficiency [24]. Effective video models incorporate temporal layers within U-Net architectures, including temporal convolution layers and temporal attention mechanisms to facilitate dynamic temporal data assimilation [25]. The integration of spatial and temporal processing remains crucial for maintaining coherent motion patterns across video sequences [26].
@@ 2.2 Linear Attention Mechanisms in Vision
The computational limitations of traditional attention mechanisms have sparked interest in linear attention alternatives that maintain performance while reducing complexity [27, 22]. Recent advances in natural language processing have introduced models with linear feature aggregation mechanisms, such as RWKV and Mamba [4, 5], which emerge as efficient solutions for processing lengthy sequences while maintaining transformer-like capabilities. Complementary work has explored linear attention from kernel-based and low-rank approximation perspectives. Performer [28] employs FAVOR+ kernel projections to approximate softmax attention with linear complexity, enabling efficient global receptive fields in vision tasks. Linformer [29] reduces attention cost via low-rank projections of key and value matrices, demonstrating that self-attention can be inherently low-dimensional in many vision backbones. RetNet [30] further advances this direction by unifying recurrence and attention through retention mechanisms, offering a scalable alternative for long-context modeling in vision and multimodal workloads. In the vision domain, these mechanisms have increasingly been adapted to high-resolution and generative modeling tasks. Diffusion-RWKV [31] shows that RWKV-based architectures can efficiently process patchified inputs while significantly reducing spatial aggregation costs, achieving competitive performance without windowing or cached-group operations. VRWKV [8] extends these principles to visual modeling, combining linear complexity with strong global context processing. Its bidirectional Bi-WKV mechanism supports comprehensive spatial–temporal reasoning, making it particularly effective for tasks such as video editing. Together, these works highlight the growing viability of linear attention as an efficient and scalable alternative to transformer-based vision architectures.
@@ 3 Background
This section provides the necessary foundations that are used in our approach. First, we discuss latent variable refinement techniques that improve the efficiency and quality trade-off. We then review diffusion models in the context of video generation and highlight their extension from image-based frameworks to spatio-temporal domains, followed by a detailed overview of the VRWKV architecture, which offers a linear complexity alternative to traditional attention mechanisms. These components form the conceptual backbone for our proposed method.
@@ 3.1 Diffusion Models and Latent Variable Refinement
Diffusion models represent [39] a powerful class of generative models that operate through a progressive noise addition and removal process. Starting with an original image $x_{0}$, the forward diffusion process gradually corrupts the data by adding Gaussian noise over $T$ timesteps. At each step $t$, the noisy image $x_{t}$ is obtained by:
@@
The reverse process learns to denoise these corrupted images, effectively modeling the data distribution. For conditional generation tasks, we can extend this framework to model the conditional distribution $p_{\theta}(x_{t}|c,x_{0})$, where $c$ represents conditioning information such as text prompts or class labels. The conditional model learns the following distribution:
@@
where $x_{T}\sim N(0,I)$ represents pure noise. The model parameters $\theta$ are optimized by maximizing the variational lower bound of the log-likelihood:
$\mathcal{L}=\mathbb{E}_{x_{0},c,t,\epsilon}\left[\|\epsilon-\epsilon_{\theta}(x_{t},t,c)\|^{2}\right].$ (4)
The network $\epsilon_{\theta}$ learns to predict the noise added in each time step given the noisy image $x_{t}$, the time step $t$, and the conditioning information $c$.
To improve computational efficiency, latent diffusion models perform the diffusion process in a compressed latent space rather than directly in the pixel space. This approach utilizes a pre-trained autoencoder consisting of an encoder $E$ that maps images to latent representations $z=E(x)$ and a decoder $D$ that reconstructs images from latents.
The diffusion process then operates in this latent space, learning the distribution of encoded representations. This approach significantly reduces memory and computational costs while maintaining high-quality generation capabilities [40]. During inference, the model generates latent representations through the reverse diffusion process, which are then decoded back to pixel space using the pretrained decoder.
@@ 3.2 Diffusion Models for Video Processing
Beyond their success in image generation, diffusion models have recently been extended to video synthesis, where they enable high-quality spatio-temporal generation and editing [41, 42, 43]. These models operate by learning to reverse a gradual noise addition process, where data is progressively corrupted through a forward Markov chain and then restored through a learned reverse process.
The transition from image to video generation involves inflating 2D models to handle spatiotemporal dimensions, typically through pseudo-3D convolutions where 3×3 kernels are replaced by 1×3×3 kernels, and temporal attention layers are added to transformer blocks [23, 44]. However, this extension introduces significant computational challenges due to the quadratic complexity of attention mechanisms when processing multiple frames simultaneously.
@@ 3.3 VRWKV Architecture
VRWKV consists of three main components: Receptance $(R)$, Key $(K)$, and Value $(V)$ transformations, combined with a time-mixing mechanism specifically adapted for vision tasks. For a given input sequence of video frames $X_{t}\in\mathbb{R}^{f\times c\times h\times w}$ consisting of $f$ frames of size $h\times w$ and the number of channels $c$, the VRWKV operations are defined as:
@@
where $W_{R}$, $W_{K}$, $W_{V}$ are learnable projection matrices, $\mu_{R}$, $\mu_{K}$, $\mu_{V}$ are learnable interpolation factors and $\otimes$ denotes a spatiotemporal convolution, defined as a standard $2D$ spatial convolution applied on a temporally interpolated frame representation $\mu\odot X_{t}+(1-\mu)\odot X_{t-1}$. This differs from standard 3D convolution since it only considers the current and previous frame, rather than a full temporal kernel.
@@
$wkv_{t}=\operatorname{Bi-WKV}(K,V)_{t}=\frac{\sum_{i=0,i\neq t}^{T-1}e^{-(|t-i|-1)/T\cdot w+k_{i}}v_{i}+e^{u+k_{t}}v_{t}}{\sum_{i=0,i\neq t}^{T-1}e^{-(|t-i|-1)/T\cdot w+k_{i}}+e^{u+k_{t}}}$ (8)
where $T$ represents the total number of tokens, $w$ and $u$ are two $d$-dimensional learnable vectors that represent channel-wise spatial decay and the bonus indicating the current token, respectively. $k_{t}$ and $v_{t}$ denote $t$-th feature of $K$ and $V$.
@@
$Y_{t}=X_{t}+O_{t}+CM_{t}.$ (13)
@@
Figure 2: Pipeline of VRWKV-Editor: Given a text–video pair (e.g., “A man with a backpack hikes on a rocky terrain”) as input, our method leverages pretrained text-to-image diffusion models for text-to-video generation. The input video is first encoded into a discrete latent space, after which our U-Net architecture predicts the injected noise (a detailed illustration is depicted in Figure 3). During inference, a novel video is synthesized by inverting the discrete noise from the input video, guided by an edited prompt (e.g., “An astronaut with a jetpack floats above a Martian landscape, with red rocky terrains and tall”).
@@ 4 Method
This section presents our contribution. We first formulate the problem and highlight the computational efficiency issues in existing architectures, then detail the proposed method.
@@ 4.1 Problem Formulation
Given a video sequence $X=\{X_{i}\}_{1}^{n}$ with $n$ frames, and a text prompt $P$ describing $X$, our goal is to generate an edited video $X^{{}^{\prime}}=\{X_{i}^{{}^{\prime}}\}_{1}^{n}$ that aligns with new prompts $P^{{}^{\prime}}$ provided by the user, where $X_{i}^{{}^{\prime}}\in\mathbb{R}^{C\times H\times W}$ is a frame with C channels, height H and width W.
@@ 4.2 Computational Efficiency Challenges
where Q, K and V represents the Querie, Key and Value, respectively. Although effective in capturing long-range dependencies, this operation scales poorly with sequence length.
@@
4.3 The proposed architecture: VRWKV-Editor
@@
Tune-A-Video [11] reduced the high computational complexity of spatio-temporal attention (ST-Attn) in generating videos with increasing frames by introducing a novel sparse version of causal attention mechanism, where the attention matrix is computed between frame $X_{i}$ and two previous frames $X_{1}$ and $X_{i-1}$, reducing the complexity from $O((dT)^{2})$ to $O(2dT^{2})$. However, the quadratic complexity still remains a bottleneck for long sequences of video frames, limiting their ability to efficiently process high-resolution images and lengthy sequences, posing a significant barrier to their broader application. To overcome this issue, we propose a linear attention mechanism based on the VRWKV [8] framework, where we replace the traditional quadratic attention computation with a receptance-weighted key-value operation that scales linearly with sequence length, reducing the complexity from $O(2dT^{2})$ to $O(2dT)$. Before aggregation, VRWKV employs a lightweight mixing mechanism to capture local temporal dynamics without heavy convolution. For a given input sequence of video frames $X_{t}$, we compute the Receptance $(R_{t})$, Key $(K_{t})$, and Value $(V_{t})$ vectors using a learnable interpolation with the previous time-step state $X_{t-1}$ as defined in Eq. 7. In the standard Transformer, the softmax prevents the decomposition of the attention matrix. VRWKV replaces this with a channel-wise exponential decay (see Eq.8). By defining the accumulated states $A_{t}$ and $B_{t}$:
$A_{t}=\sum_{i=0}^{t}e^{-(t-i)w+k_{i}}v_{i},\quad B_{t}=\sum_{i=0}^{t}e^{-(t-i)w+k_{i}}$ (17)
We can express the state update recursively:
$A_{t}=e^{-w}\odot A_{t-1}+e^{k_{t}}\odot v_{t}$ (18)
$B_{t}=e^{-w}\odot B_{t-1}+e^{k_{t}}$
This recursive formulation demonstrates that calculating the output for step $t$ depends only on the fixed-size hidden states $A_{t-1}$ and $B_{t-1}$, not on the entire history of previous frames. Consequently, the complexity collapses from $O(2dT^{2})$ to $O(2dT)$.
Our approach leverages the VRWKV’s inherent ability to capture long-range dependencies through its recurrent formulation while maintaining spatial awareness through learnable shift operations and channel mixing.
@@
An overview of our approach is depicted in Figure 2. Given a text-video pair as input, our model leverages the pretrained T2I diffusion models for T2V generation by encoding the input video into a discrete space, then we predict the added noise using our U-Net architecture enhanced with VRWKV modules (Figure 3). During inference, we sample a novel video from the discrete noise inverted from the input video, guided by an edited prompt.
4.4 3D-VRWKV Module
Figure 3: Architecture of the U-Net backbone employed in VRWKV-Editor. The design incorporates VRWKV modules within the skip connections, enabling efficient long-range dependency modeling while preserving fine-grained spatial information across down-sampling and up-sampling layers.
Figure 4: VRWKV [8] module employed on VRWKV-Editor. The 3D-VRWKV architecture includes identical VRWKV encoder layer, an average pooling layer, and a linear prediction head. Token Shift denotes the quad-directional shift method tailed for vision tasks.
@@
During our experiments, we observed that traditional CNN-based spatiotemporal modules failed to capture complex spatiotemporal dependencies, while adding more cross-attentions or temporal attentions [47, 48] only increased the complexity of the model without improving the quality of edited videos or maintaining their consistency. Inspired by the RWKV architecture’s ability to combine the strengths of RNNs and Transformers, we introduce 3D-VRWKV (Figure 4), a variant of VRWKV module specifically designed for video spatiotemporal modeling. The 3D-VRWKV module consists of two sequential sub-modules: spatiotemporal mixing and channel mixing, each operating with linear complexity.
Spatiotemporal Mixing with Bidirectional Processing: The spatiotemporal mixing module processes input video features $X$ through temporal interpolation, spatial projection and Bidirectional Weighted Key-Value Attention.
To enable the model to learn smooth temporal transitions, we first mix each frame’s features with those of its predecessor:
$X^{mix}_{t}=\mu\odot X_{t}+(1-\mu)\odot X_{t-1},\quad t=1,...,T$ (19)
where $\mu\in\mathbb{R}^{1\times 1\times 1\times 1\times C}$ is a learnable channel-wise interpolation parameter.
For the spatial projection, the mixed features are projected to receptance, key, and value spaces through learned transformations:
$R_{s}$ $=W_{R}\star X_{t}^{\text{mix}}\in\mathbb{R}^{B\times H\times W\times d}$ (20)
$K_{s}$ $=W_{K}\star X_{t}^{\text{mix}}\in\mathbb{R}^{B\times H\times W\times d}$ (21)
$V_{s}$ $=W_{V}\star{X}_{t}^{\text{mix}}\in\mathbb{R}^{B\times H\times W\times d}$ (22)
where $\star$ denotes spatial convolution (implemented as 1×1 convolutions for computational efficiency), and $d=H_{\text{heads}}\times d_{\text{head}}$ is the total embedding dimension. We reshape these projections to separate the head dimension: $R_{s},K_{s},V_{s}\in\mathbb{R}^{B\times T\times N\times H_{\text{heads}}\times d_{\text{head}}}$ where $N=H\times W$ is the total number of spatial tokens.
Channel Mixing: After spatiotemporal mixing, we apply a channel mixing module to enable communication across feature channels, which is crucial for modeling correlations between different semantic attributes such as object identity, color, and texture. Similar to the spatiotemporal mixing stage, we begin by creating temporally interpolated channel features through:
$Y^{cmix}_{t}=\mu_{c}\odot Y^{spatial}_{t}+(1-\mu_{c})\odot Y^{spatial}_{t-1}$ (23)
where $\mu_{c}\in\mathbb{R}^{1\times 1\times 1\times 1\times C}$ is a separate learnable parameter for channel mixing that allows the model to adaptively control temporal smoothness in the channel domain. These interpolated features are then projected to receptance and key spaces through learned transformations:
$R_{t}^{c}$ $=W_{R}^{c}\star Y_{t}^{\text{cmix}}$ (24)
${K}_{t}^{c}$ $={W}_{K}^{c}\star{Y}_{t}^{\text{cmix}}$ (25)
where $\star$ denotes spatial convolution implemented as $1\times 1$ convolutions.
The channel mixing output is then computed through a squared ReLU activation that provides enhanced non-linearity for modeling complex channel interactions:
$\text{CM}_{t}=\sigma(R^{c}_{t})\odot\max(K^{c}_{t}\cdot W_{c},0)^{2}$ (26)
where $W_{c}$ is an additional learnable transformation matrix and the squared activation $\max(\cdot,0)^{2}$ provides stronger non-linearity than standard ReLU. The final output combines both spatiotemporal and channel mixing through residual connections:
$Z_{t}=X_{t}+Y^{\text{spatial}}_{t}+\text{CM}_{t}$ (27)
The 3D-VRWKV module addresses key limitations of traditional attention mechanisms in video editing. Unlike quadratic attention mechanisms, 3D-VRWKV operates with linear complexity $O(d)$, making it more efficient for processing long video sequences (See Table 2). The module maintains a recurrent state that captures temporal dependencies across frames while preserving spatial relationships. Moreover, the weighted key-value mechanism allows for dynamic focus on relevant spatiotemporal features.
@@
Figure 5: Qualitative comparison between CAMEL [49], CCEdit [50], ControlVideo [20], Vid2Vid-zero [10], Swin-Editor [51] and our method. The first row presents the input video. The figure highlights two different scenarios: Object editing and object-background editing . The results showcase the superiority of our framework in editing videos across various scenarios, producing consistent and high quality outputs.